# Advancing MAPF towards the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)

### v2

@@ III-D Performance Metrics
IV Using SMART
Users can evaluate their solutions using SMART in two ways: by uploading their MAPF problem and solution through a web interface or by deploying SMART locally.
Figure 5: The SMART web interface visualizing a multi-robot simulation on a grid map. In the viewport, users interact with individual robots to inspect paths and inter-robot constraints. On the side panel, users can configure the environment, control playback, and see execution details.
@@ IV Using SMART
SMART, with ARGoS3 and the 3D visualizer, is directly available as software-as-a-service [^2] The link will be made available here upon acceptance.. Specifically, we provide an intuitive web app for running and inspecting MAPF simulations. Complete with a beautiful 3D-visualizer built with React and Three.js, it allows researchers to instantly understand how algorithms perform with physical robots (Fig. 5). Users configure the environment by specifying a map and scenario file in the MovingAI Benchmark format. Within the online interface, after users upload the files and select the desired kinodynamic settings for the robots, and initiate the simulation with a single click. After starting the simulation, they can control the simulation time step and adjust the playback speed. Users may select robots to view complete paths, execution status, and Type-2 edges between robots.
@@ V System Evaluation
In all experiments, the simulation update period is set to 0.1 simulation seconds for both simulators. A simulation update refers to the periodic recalculation of the status of robots and the environment based on the input commands. In ARGoS3, each grid cell represents 1 meter. Each foot-bot robot is subject to a speed limit of [0, 5] m/s, an acceleration limit of [-0.4, 0.4] m/s2, and an angular velocity limit of 30/s. In Isaac Sim, each grid cell represents 0.5 meters. Each JetBot is constrained to a speed limit of [0, 2.5] m/s, an acceleration limit of [-0.2, 0.2] m/s2, and the same angular velocity limit of 30/s.
We evaluated our system using MAPF plans generated by three different methods: CBS [12], PBS [7], and MAPF-LNS2 [9]. These plans were created using four different robot models: the standard MAPF model [1], the $k$-robust delay model [19], MAPF with rotation [33], and MAPF with kinodynamics [21]. Across all combinations, our system achieved a 100% execution success rate. In the remainder of this section, we present results based on MAPF plans generated by MAPF-LNS2, as it demonstrated the highest scalability among the algorithms evaluated.
@@ V-A Scalability and Runtime Performance
This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the simulation time to the real-world time. A higher simulation speed means that more simulation updates can be finished within the same unit of time, indicating better scalability. For example, if the simulation speed is 10, the simulation world runs 10 times faster than in real time. MAPF plans are generated using MAPF-LNS2333Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 [9] with a progressively increasing number of robots and a time limit of 60 seconds. For each map, we randomly select 5 scenarios from the MovingAI Benchmark to generate plans, which are then executed in SMART without visualization. In ARGoS3, the maximum number of robots tested corresponds to the limit imposed by the MAPF instances from the MovingAI Benchmark for all maps except sortation-center, which is set to 2,000 robots. In contrast, for Isaac Sim, the upper bound on the number of robots is determined by the point at which the simulation speed drops below real-time (i.e., simulation speed << 1).
@@
TABLE I: Replicability using different numbers of robots.
Figure 7: Simulation speed on all environments.
As shown in Fig. 7, SMART supports simulations with up to 2,000 robots. The more robots and the larger the map, the smaller the simulation speed. Nevertheless, for ARGoS3, SMART maintains a simulation speed of around 10.0 with 1,000 robots in warehouse. Even in the larger sortation-center, the simulation speed is consistently greater than 1.0. In contrast, Isaac Sim provides the capability to model more complex real-world factors while also producing photorealistic visualizations, but it is significantly less scalable, particularly in large environments. It is therefore more suitable for experiments with a smaller number of agents.
@@ V-B Replicability
We evaluate the consistency of SMART by examining whether the same MAPF plans yield the same outcomes across multiple runs. Due to the realistic setting we used, uncertainty may arise from factors such as network communication delays or controller variability. Using the same warehouse map and MAPF plans as in the previous section, we execute SMART 10 times for each MAPF plan. For each run, we record both the average and the maximum execution time. As shown in Table I, SMART demonstrates strong replicability across runs, with minor execution time variations. While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
@@ V-C CPU and Memory Usage
We evaluate the CPU and memory usage of SMART on both machines (1) and (2). We use machine (2) to demonstrate that SMART can run on resource-constrained machines. To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark [1]. All simulations are parallelized on all CPU cores on both machines. We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots. Table II shows the results. When using ARGoS3 as the simulator, we observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops. The simulation speed in machine (2) is slower than in machine (1) due to an inferior processor. In contrast, simulations with Isaac Sim result in significantly higher CPU and memory usage, indicating that Isaac Sim is better suited for scenarios with fewer agents but more powerful hardware.
@@
TABLE II: Comparison of CPU usage, memory consumption, and simulation speed for ARGoS3 and Isaac Sim across two machines and various robot counts. Since Isaac Sim only supports recent GPUs with ray tracing capabilities, we conducted experiments with Isaac Sim only on Machine (1).
@@ VI Conclusion
We introduce SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings. SMART bridges the gap between methods that use simplified MAPF models and their real-world deployments. SMART is compatible with existing 2-D grid-based benchmarks, making it straightforward for researchers to evaluate their methods in more realistic environments. We integrate SMART with two simulators, ARGoS3 and Isaac Sim. Our experiments show that SMART can handle thousands of robots while maintaining high simulation speeds when using ARGoS3. Future work includes extending the framework to address lifelong MAPF scenarios and integrating it with other execution frameworks [34, 35, 36].

### v3

@@ III-D Performance Metrics
IV System Usage
Users can evaluate their solutions using SMART in two ways: by uploading their MAPF problem instance and plan through a web interface or by deploying SMART locally.
Figure 5: The SMART web interface visualizing a multi-robot simulation on a grid map. In the viewport, users interact with individual robots to inspect paths and inter-robot constraints. On the side panel, users can configure the environment, robot kinematic limits, control playback, and see execution details.
@@ IV System Usage
SMART, with ARGoS3 and the 3D visualizer, is directly available as software-as-a-service. Specifically, we provide an intuitive web app for running and inspecting MAPF simulations. It features a 3D visualizer built with React and Three.js, allowing researchers to better understand how algorithms perform on physical robots (Fig. 5). Users configure the environment by specifying a map and scenario file in the MovingAI Benchmark format. Within the online interface, after users upload the files and select the desired kinodynamic settings for the robots, users can initiate the simulation with a single click. After starting the simulation, they can control the simulation time step and adjust the playback speed. Users may select robots to view full paths, execution status, and Type-2 edges between robots.
@@ V System Evaluation
In all experiments, the simulation update period is set to 0.1 simulation seconds for both simulators. A simulation update refers to the periodic recalculation of the status of robots and the environment based on the input commands. In ARGoS3, each grid cell represents 1 meter. Each foot-bot robot is subject to speed limits of [0, 5] m/s, acceleration limits of [-0.4, 0.4] m/s2, and absolute angular velocity limits of 30/s. In Isaac Sim, each grid cell represents 0.5 meters. Each JetBot is limited to a speed limit of [0, 2.5] m/s, an acceleration limit of [-0.2, 0.2] m/s2, and the angular velocity limit of 30/s.
We evaluate SMART using MAPF plans generated by MAPF-LNS2 [14], which provides strong scalability and is representative of state-of-the-art MAPF solvers. Unless otherwise specified, all results in this section are based on MAPF-LNS2. To verify generality, we additionally conducted experiments using plans generated by other MAPF planners, including CBS [28] and PBS [16]. These plans were created using four different robot models: the standard MAPF model [29], the $k$-robust delay model [1], MAPF with rotation [36], and MAPF with kinodynamics [34]. Across all tested settings, SMART achieved a 100% execution success rate, demonstrating its compatibility with different MAPF planners.
@@ V-A Scalability and Runtime Performance
This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the simulation time to the real-world time. A higher simulation speed means that more simulation updates can be finished within the same unit of time, indicating better scalability. For example, if the simulation speed is 10, the simulation world runs 10 times faster than in real time. MAPF plans are generated using MAPF-LNS2222Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 [14] with a progressively increasing number of robots and a time limit of 60 seconds. For each map, we randomly select 5 scenarios from the MovingAI Benchmark to generate plans, which are then executed in SMART without visualization. In ARGoS3, the maximum number of robots tested corresponds to the limit imposed by the MAPF instances from the MovingAI Benchmark for all maps except sortation-center, which is set to 2,000 robots. In contrast, for Isaac Sim, the upper bound on the number of robots is determined by the point at which the simulation speed drops below real-time (i.e., simulation speed << 1).
Figure 7: Simulation speed in all environments.
TABLE II: Replicability using different numbers of robots.
@@
As shown in Fig. 7, SMART supports simulations with up to 2,000 robots. As the number of robots increases and the map becomes larger, the simulation speed decreases. Nevertheless, for ARGoS3, SMART maintains a simulation speed of around 10.0 with 1,000 robots in warehouse. Even in the larger sortation-center, the simulation speed is consistently greater than 1.0. In contrast, Isaac Sim provides the capability to model more complex real-world factors while also producing photorealistic visualizations, but it is significantly less scalable, particularly in large environments. It is therefore more suitable for experiments with a smaller number of robots.
@@ V-B Replicability
We evaluate the consistency of SMART by examining whether the same MAPF plans yield the same outcomes across multiple runs. Due to the realistic setting we used, uncertainty may arise from factors such as network communication delays or controller variability. Using the same warehouse map and MAPF plans as in the previous section, we execute SMART 10 times for each MAPF plan. For each run, we record both the average and the maximum execution time. As shown in Table II, SMART demonstrates strong replicability across runs, with minor execution time variations. While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
@@ V-C CPU and Memory Usage
We evaluate the CPU and memory usage of SMART on both machines (1) and (2). We use machine (2) to demonstrate that SMART can run on resource-constrained machines. To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MovingAI Benchmark [29]. All simulations are parallelized on all CPU cores on both machines. We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots. Table III shows the results. When using ARGoS3 as the simulator, we observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops. The simulation speed in machine (2) is slower than in machine (1) due to an inferior processor. In contrast, simulations with Isaac Sim result in significantly higher CPU and memory usage, indicating that Isaac Sim is better suited for scenarios with fewer robots but more powerful hardware.
TABLE III: Comparison of CPU usage, memory consumption, and simulation speed for ARGoS3 and Isaac Sim across two machines and various robot counts. Since Isaac Sim only supports recent GPUs with ray tracing capabilities, we conduct experiments with Isaac Sim only on Machine (1).
@@ VI Conclusion
We introduce SMART, a scalable testbed for evaluating MAPF planners under realistic settings. SMART bridges the gap between methods that use simplified MAPF models and their real-world deployments. SMART is compatible with existing 2-D grid-based benchmarks, making it straightforward for researchers to evaluate their methods in more realistic environments. We integrate SMART with two simulators, ARGoS3 and Isaac Sim. Our experiments show that SMART can handle thousands of robots while maintaining high simulation speeds when using ARGoS3. SMART’s modular, planner-agnostic design supports diverse application-driven use cases across academic and industrial settings. Academic researchers can easily benchmark different algorithms without custom integration, while industrial users can evaluate algorithms in their specific warehouse or manufacturing environments without requiring specialized MAPF expertise. Future work includes: (1) extending to lifelong MAPF [17] where tasks arrive continuously and require dynamic replanning, (2) integrating sophisticated dynamic obstacle models for human-robot interaction scenarios, (3) supporting alternative execution frameworks [6, 2, 7].
Acknowledgments
The research was supported by the National Science Foundation (NSF) under grant numbers #2328671 and #2441629, as well as a gift from Amazon. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government.